OpenStudy (anonymous):

Find the slope of the line that contains the points, (1, 2) and (2, 5). Part b.) Using the slope that you found in part a and one of the given points, write the equation of the line. Please write your final answer in slope-intercept form.

OpenStudy (radar):

Part a) find slope. m=slope \[m=\Delta y/\Delta x\] \[\Delta y=(2-5)=-3\]\[\Delta x=1-2=-1\]\[m=-3/-1=3\]

OpenStudy (radar):

Slope is 3. Do you need help for the b.) part ?

OpenStudy (anonymous):

yess please

OpenStudy (radar):

The slope is equal to (y-y1)/(x-x1) where (x1, y1) is one of your points, I am choosing the first point (1, 2) Substituting 3 for m (slope) 3=(y-2)/(x-1) Cross multiply and place result in the slope intercept form 3x-3=y-2 y=3x-1 Thats it the coefficient of x is the slope and the constant (-1) is the intercept.

OpenStudy (radar):

Did you follow with understanding?

OpenStudy (anonymous):

yes i did is that all of part b?

OpenStudy (radar):

yes, part b just wanted you to "write the equation of the line" "in slope intercept form" that would be y=3x-1.

OpenStudy (anonymous):

thank u soo much

OpenStudy (anonymous):

Part a.) Write your own function in function notation. For example: f(x) = x + 1. Using your function, find f(2). (5 points) Part b.) Explain in your own words how to find the value of f(2) when you are given a function f(x). (5 points)

OpenStudy (anonymous):

can u help